OpenStudy (anonymous):

Solve the quadratic equation using the square root property: (x+4)^2 = -81

OpenStudy (anonymous):

So I set it to x^2 = -85 and now I'm confused because it won't take my answer of sqrt85 or -sqrt 85

OpenStudy (owlcoffee):

Oh no, my friend. Uou can't take out the 4 just like that, you have to expand that little (x+4)^2

OpenStudy (anonymous):

so should I make it x^2 +16 and then x^2= -97???

OpenStudy (owlcoffee):

oh, no no. Let's do this: \[(a+b)^{2}\] This does mean : \[(a+b)(a+b)\] Now, let's apply distributive there: \[a ^{2}+ab+ba+b ^{2}\] By commutative I know that ab and ba are equal: then: \[a ^{2}+2ab+b ^{2}\] That would be the formula you have to apply, Look at your problem, a=x and b=4. Can you do that?

OpenStudy (anonymous):

x^2 +(2x*8) + 16?

OpenStudy (anonymous):

x^2+16x+16? So then what do I do with that?

OpenStudy (owlcoffee):

solve it for x using sqrt properties :)

OpenStudy (anonymous):

So then I'd make it x^2 +16x= -97? Oof, I feel like I just keep missing osmething here

OpenStudy (owlcoffee):

Hmm... Let's try something. pass that -81 to the other side as negative, and transform it to a 9^2 Use some exponential properties on those.

OpenStudy (anonymous):

sqrtx^2 + sqrt26x + sqrt16 + 9^2 = 0?

OpenStudy (anonymous):

I mean to erase the sqrts on those, oops

OpenStudy (anonymous):

does x = -13